OpenStudy (anonymous):

Find the zeros of the function h(x) = x2 +4x by factoring. What is the smallest zero of the function?

OpenStudy (lgbasallote):

@taliii what is the factored form of x^2 + 4x?

OpenStudy (anonymous):

\[\huge \color{green}{\textbf{Welcome to Openstudy}}\]

OpenStudy (anonymous):

Will you write it today?? Ha ha ha..

OpenStudy (lgbasallote):

here's my own personal taste: \(\Huge \color{red}{\mathbb{WELCOME}}\) \(\huge \color{green}{\mathbf{TO}}\) \(\Huge \color{fuchsia}{\mathcal{OPENSTUDY!!!}}\)

OpenStudy (lgbasallote):

heh i'm not as fast as i used to

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 + 4x = 0}\) \( \color{Black}{\Rightarrow x(x + 4) = 0}\) The smaller solution will be the answer.